# Z-Image: An Efficient Image Generation Foundation Model with Single-Stream Diffusion Transformer

### v2

@@ 5.1.1 Elo-based Human Preference Evaluation
To rigorously benchmark Z-Image-Turbo’s capabilities against the competitive landscape of generative models, we participated in Alibaba AI Arena [^2] https://aiarena.alibaba-inc.com/corpora/arena/leaderboard?arenaType=T2I, a public-facing, independent benchmarking platform powered by large-scale human judgment. Unlike automated metrics that frequently misalign with human perception, the Arena provides dynamic and unbiased model rankings based on thousands of pairwise comparisons, making it an ideal venue for objective performance assessment.
In the evaluation, Z-Image-Turbo – our high-efficiency diffusion architecture with 6B parameters and low inference cost of 8 NFEs – competed against 8 leading models, including top-tier closed-source systems: Imagen 4 Ultra Preview 0606 google2025imagen4, gemini-2.5-flash-image-preview google2025gemini2.5flashimage (both Google), Seedream 4.0 (ByteDance) seedream2025seedream, GPT Image 1 [High] (OpenAI) openai2025gptimage, FLUX.1 Kontext [Pro] (Black Forest Labs) labs2025flux, and Ideogram 3.0 ideogramV3, along with the open-source baseline Qwen-Image qwenimage.
As shown in Table 3, Z-Image-Turbo attained an Elo score of 1025, achieving 4th place globally and ranking 1st among open models on the leaderboard [^3] Snapshot taken on November 26, 2025. Despite its efficient design, it outperforms multiple industry-leading closed-source models and surpasses Qwen-Image while operating at significantly lower computational cost. With a 45% win rate across all matchups, Z-Image-Turbo demonstrates strong competitiveness against the entire field, including top-tier proprietary systems. This high level of performance positions it firmly within the elite tier of generative models, affirming its status as one of the most advanced open-source systems available.
These results establish Z-Image-Turbo as one of the leading open text-to-image models in terms of both quality and efficiency. More than a high-performing generator, it represents a new baseline for efficiency-oriented architecture design, demonstrating that compact models can achieve elite-level performance without compromising usability. This combination of speed, fidelity, and openness enables deployment in resource-constrained environments, interactive applications, and community-driven innovation.
@@ 5.3.1 Superior Photorealistic Generation
As shown in Figure 16 and 17, Z-Image-Turbo shows excellent character close-up generation (e.g., the skin details on a man’s face and a girl’s tears). When asked to generate multi-expression portraits of one person (Figure 18), Z-Image-Turbo can produce images that are more aesthetically pleasing and have more realistic expressions, while Qwen-Image, HunyuanImage3.0, FLUX 2 dev, and Seedream 4.0 would sometimes generate exaggerated and unrealistic expressions, thus lacking authenticity and beauty. Moreover, when generating a scene captured by a mobile phone (Figure 19 and 20), Z-Image-Turbo shows strong performance in the authenticity of both the person and the background, as well as the aesthetic appeal of layout and posture. while Qwen-Image, HunyuanImage3.0, and FLUX 2 dev would generate unrealistic things (e.g., clothes that remain completely unsoaked in the heavy rain).
@@ 5.3.4 Enhanced Reasoning Capacity and World Knowledge through Prompt Enhancer
As demonstrated in Figure 15 and Figures 26-27, our prompt enhancer leverages a structured reasoning chain – comprising core subject analysis, problem solving/world knowledge injection, aesthetic enhancement, and comprehensive description – to equip the model with logical reasoning and world knowledge capabilities. This allows the model to handle diverse tasks, ranging from solving complex logical puzzles (e.g., the chicken-and-rabbit problem) and interpreting user intent (e.g., visualizing classical poetry or inferring scenes from coordinates) to performing text rendering and question answering.
In the context of image editing, prompt enhancer is also crucial for addressing ambiguous or unclear intentions, as well as for injecting world knowledge and enabling reasoning, similar to how it functions in text-to-image generation, as shown in Figure 28-29. For example, in Figure 29, the wrong dish is made because of a lack of reasoning about the relationship between the ingredients and the dish, while prompt enhancer can make up for this.

### v3

@@ 5.1.1 Elo-based Human Preference Evaluation
Figure 16: Z-Image-Turbo secured the 8th position among all evaluated models in the Text-to-Image Elo rankings provided by the Artificial Analysis AI Arena.
Figure 17: Z-Image-Turbo ranks first among open-source models in the Text-to-Image Elo rankings from the Artificial Analysis AI Arena.
To rigorously benchmark Z-Image-Turbo’s capabilities against the competitive landscape of generative models, we participated in Artificial Analysis Image Arena [^2] https://artificialanalysis.ai/image/leaderboard/text-to-image and Alibaba AI Arena [^3] https://aiarena.alibaba-inc.com/corpora/arena/leaderboard?arenaType=T2I, two public-facing, independent benchmarking platforms powered by Elo-based large-scale human judgment. Unlike automated metrics that frequently misalign with human perception, the Elo-based evaluation provides dynamic and unbiased model rankings based on thousands of pairwise comparisons, making it an ideal venue for objective performance assessment.
In the evaluation, Z-Image-Turbo – our high-efficiency diffusion architecture with 6B parameters and a low inference cost of 8 NFEs – was benchmarked against other leading models. These included top-tier closed-source systems such as Nano Banana Pro nanopro, Imagen 4 Ultra Preview 0606 google2025imagen4 (Google), Flux.2 (Black Forest Labs) labs2025flux, Seedream (ByteDance) seedream2025seedream, GPT Image 1 [High] (OpenAI) openai2025gptimage, and FLUX.1 Kontext [Pro] (Black Forest Labs) labs2025flux, as well as open-source baselines like HunyuanImage 3.0 cao2025hunyuanimage and Qwen-Image qwenimage.
According to the Artificial Analysis leaderboard in Figure 16 and Figure 17, Z-Image-Turbo achieved an Elo score of 1,161, ranking 8th overall. This result is notable for two primary reasons. Firstly, Z-Image-Turbo is established as the top-ranked open-source model on this highly competitive platform, with performance comparable to that of leading proprietary systems like Google’s Imagen 4 and ByteDance’s Seedream. Secondly, the model exhibits exceptional efficiency: among the top ten models, it features not only the smallest parameter count (6B) but also the lowest inference cost (\$5.0 per 1,000 images). This unique combination of high-quality output and low computational overhead is a defining characteristic of our architecture. These findings are further corroborated by its performance on the Alibaba AI Arena (see Table 3), where Z-Image-Turbo secures a higher rank of 4th overall while again leading the open-source category.
In summary, these results establish Z-Image-Turbo as one of the leading open text-to-image models in terms of both quality and efficiency. More than a high-performing generator, it represents a new baseline for efficiency-oriented architecture design, demonstrating that compact models can achieve elite-level performance without compromising usability. This combination of speed, fidelity, and openness enables deployment in resource-constrained environments, interactive applications, and community-driven innovation.
@@ 5.3.1 Superior Photorealistic Generation
As shown in Figure 18 and 19, Z-Image-Turbo shows excellent character close-up generation (e.g., the skin details on a man’s face and a girl’s tears). When asked to generate multi-expression portraits of one person (Figure 20), Z-Image-Turbo can produce images that are more aesthetically pleasing and have more realistic expressions, while Qwen-Image, HunyuanImage3.0, FLUX 2 dev, and Seedream 4.0 would sometimes generate exaggerated and unrealistic expressions, thus lacking authenticity and beauty. Moreover, when generating a scene captured by a mobile phone (Figure 21 and 22), Z-Image-Turbo shows strong performance in the authenticity of both the person and the background, as well as the aesthetic appeal of layout and posture. while Qwen-Image, HunyuanImage3.0, and FLUX 2 dev would generate unrealistic things (e.g., clothes that remain completely unsoaked in the heavy rain).
@@ 5.3.4 Enhanced Reasoning Capacity and World Knowledge through Prompt Enhancer
As demonstrated in Figure 15 and Figures 28-29, our prompt enhancer leverages a structured reasoning chain – comprising core subject analysis, problem solving/world knowledge injection, aesthetic enhancement, and comprehensive description – to equip the model with logical reasoning and world knowledge capabilities. This allows the model to handle diverse tasks, ranging from solving complex logical puzzles (e.g., the chicken-and-rabbit problem) and interpreting user intent (e.g., visualizing classical poetry or inferring scenes from coordinates) to performing text rendering and question answering.
In the context of image editing, prompt enhancer is also crucial for addressing ambiguous or unclear intentions, as well as for injecting world knowledge and enabling reasoning, similar to how it functions in text-to-image generation, as shown in Figure 30-31. For example, in Figure 31, the wrong dish is made because of a lack of reasoning about the relationship between the ingredients and the dish, while prompt enhancer can make up for this.